How much energy is given off or absorbed when 26 mol of O2 is reacted?
CH4(g) + 2O2(g) → CO2(g) + 2H2O(l) ΔH = –890 kJ
1
Expert's answer
2020-01-28T06:15:32-0500

"CH_4(g) + 2O_2(g) → CO_2(g) + 2H_2O(l) ;\ ΔH = –890 kJ"


The stochiometric coefficient of oxygen in the above reaction is 2 which means when 2 moles of oxygen is reacted 890 kJ of heat is released

When one moles of oxygen is reacted then 890/2(=445 kJ) of energy is released

So, when 26 moles of oxygen is reacted then energy released = 445*26 = 11570 kJ